OpenStudy (anonymous):

(3-5)*4-(-4-5)to the second power

OpenStudy (blurbendy):

(3-5)*4-(-4-5)^2 (-4 -5)(-4-5) = 16 + 40 + 25 = 65 + 16 = 81 4(3-5) = 12 -20 = -8 -8 - 81 = -90